OpenStudy (redstealth29):

Sarah burns an average of 744 calories in 45 minutes while running. She ran at 10 miles per hour. There are 5,280 feet per a mile. How many calories did Sarah burn per foot? Round to the nearest ten-thousandth.

OpenStudy (redstealth29):

Could anyone help me with this?

OpenStudy (el_tucan):

this sounds like a proportion... let me think about this... the first thing that came to mind was R = D*T

OpenStudy (redstealth29):

I don't understand @El_Tucan

OpenStudy (el_tucan):

give me a sec... i was lost but i have a friend helping me...

OpenStudy (el_tucan):

you know that 744 calories are burned in 45 mins and also know that the rate she is moving at is 10mph. the question wants to know how many calories she burns per foot so u must convert minutes all the way to miles using information given. start with ratio given which is : 744 cal / 45 mins you want to go from minutes to hours because her speed is in mph so you multiply the ratio above by 60 min / 1 hr so that you end up with: 744 cal / 45 min X 60 min/ 1 hr the question gives you 10mph and in order to cancel the hour units must put the hours on top of the ratio so u end up with 744 cal / 45 min X 60 min/ 1 hr X 1hr/ 10 miles they want the amount of feet and ur given that 5280 is in a mile u put the mile units in the top of the ratio and end up with 744 cal / 45 min X 60 min/ 1 hr X 1hr/ 10 miles X 1 mile/ 5280 ft Now, minutes, hours, and miles units all cancel leaving you with calories and feet. Then, you multiply all the numerators and divide by the product of the denominators leaving you with: 744 cal / 45 min X 60 min/ 1 hr X 1hr/ 10 miles X 1 mile/ 5280 ft = 44,640 cals / 2,376,000 feet = .0187878787 cals per foot

OpenStudy (ineedserioushelp):

Alright so Sarah ran 7.65 miles in total, seeing as 10mph is 0.17 miles per minute, 45 * 0.17 = 7.65. 7.65 * 5280 = 40392. So she ran 40,392 feet, she burned 744 calories so: 744/40392 = 0.01841948900772430184194890077243. Rounded to the nearest ten-thousandth: 0.0184, so she burned 0.0184 calories per foot ran.

OpenStudy (redstealth29):

Thanks guys!